OpenStudy (anonymous):

*Find all solutions in the interval 0 < theta, 2pi for tan^2theta-3=0, and 2 sintheta+ sqrt3=0 . I'm so confused, HELP!

OpenStudy (freckles):

\[\tan^2(x)=3\] First take square root of both sides, you will end up with two equations to solve after that.

OpenStudy (freckles):

Then use the unit circle.

OpenStudy (anonymous):

\[\tan \theta=\sqrt{3}\]?

OpenStudy (freckles):

that is one equation

OpenStudy (freckles):

the second equation is tan(x)=-sqrt(3)

OpenStudy (freckles):

\[\tan^2(x)=3 \implies \tan(x)=\sqrt{3} \text{ or } \tan(x)=-\sqrt{3}\]

OpenStudy (freckles):

hint: recall: \[\frac{\frac{\sqrt{3}}{2}}{\frac{1}{2}}=\sqrt{3}\]

OpenStudy (anonymous):

Oh okay! So the solutions would be \[\pi/3, 2\pi/3, 4\pi/3, and 5\pi/3?\]

OpenStudy (freckles):

great

OpenStudy (freckles):

second one isolate the sin(x) part (also note: I'm using x instead of theta because it is just easier to type)

OpenStudy (anonymous):

Yes I figured :) Would the 2nd equation would be \[\sin \theta= -\sqrt{3}/2?\]

OpenStudy (freckles):

yep!

OpenStudy (anonymous):

Oh thank you sooo much!!!

OpenStudy (freckles):

you got it from there?

OpenStudy (anonymous):

I believe so... then the sin values would be ... 5pi/ 3, and 4pi/3?

OpenStudy (freckles):

right x or you know theta would be 5pi/3 or 4pi/3 good job you weren't that confused

OpenStudy (anonymous):

I know the unit circle, but these equations just look confusing.

OpenStudy (freckles):

you remember solving linear equations in algebra right?

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

like you know how to solve something like 2x+5=4 for x?

OpenStudy (freckles):

it is the same way you isolate the trig function then you solve for the variable inside the trig function

OpenStudy (anonymous):

I wish my teacher would just explain it like that. Would make it much clearer, lol

OpenStudy (freckles):

if this makes more sense to you replace tan(theta) with p so the first equation you first solve for p: p^2-3=0 p^2-3=0 add 3 on both sides p^2=3 take square root of both sides p=sqrt(3) or p=-sqrt(3) then replace p with tan(theta) tan(theta)=sqrt(3) or tan(theta)=-sqrt(3) second equation replace sin(theta) with p and solve for p: 2p+sqrt(3)=0 subtract sqrt(3) on both sides 2p=-sqrt(3) divide both sides by 2 p=-sqrt(3)/2 and so on...

OpenStudy (anonymous):

Yes, much easier! I thought maybe I had to use an identity or something

OpenStudy (freckles):

now I do admit this trig equations can get more complicated

OpenStudy (anonymous):

The next one is a little different. \[\cos (3\theta)=+\sqrt{3}/2\]

OpenStudy (freckles):

that one is also not too bad... is the interval the same for the solution?

OpenStudy (anonymous):

yes 0 to 360

OpenStudy (anonymous):

just divide both sides by 3?

OpenStudy (freckles):

\[\text{ so we want to solve for } \theta \text{ such that } 0 < \theta <360 \\ \text{ but first we will need to solve for } 3 \theta, \text{ let's call this } p \\ \text{ so we need to figure out the solution interval in terms of } p \text{ now } \\ p=3 \theta \implies \frac{p}{3}=\theta \\ \text{ so } 0 < \theta < 360 \implies 0 < \frac{p}{3} < 360 \implies 0(3) <p<360(3) \\ \implies 0<p<1080\] So you first need to solve \[\cos(p)=\frac{-\sqrt{3}}{2}\] on the interval (0,1080)

OpenStudy (freckles):

that is only 3 rotations around the circle

OpenStudy (freckles):

find the solutions in the first rotation then we will just add 360 to get the solutions in the second rotation and then we will add 720 to get the solutions in the third rotation

OpenStudy (freckles):

once we find p we will be able to find theta easily

OpenStudy (anonymous):

uhm... pi/6, and 5pi/6?

OpenStudy (freckles):

in the first rotation solutions should be 5pi/6 , 7pi/6

OpenStudy (freckles):

oops but we should probably use degrees since our interval is in terms of degrees

OpenStudy (anonymous):

oooh

OpenStudy (freckles):

so first rotation solutions should be 150 and 210

OpenStudy (anonymous):

why not 35?

OpenStudy (freckles):

35? you mean 30? we are looking for when cosine value is negative sqrt(3)/2 not positive second rotation solutions should be 150+360 and 210+360 third rotation solutions should be 150+720 and 210+720

OpenStudy (anonymous):

Why negative cosine?

OpenStudy (freckles):

doesn't your equation say: \[\cos(p)=-\frac{\sqrt{3}}{2} ?\] \[p=150,210,510,570,870,930 \\ \text{ remember } p=3 \theta \\ 3 \theta=150,210,510,570,870,930\]

OpenStudy (freckles):

err I'm blind I thought I seen a negative in front

OpenStudy (freckles):

you did have a positive

OpenStudy (anonymous):

hahah

OpenStudy (freckles):

but maybe I should finish solving this one to give you an idea for the positive version of sqrt(3)/2 for cos

OpenStudy (anonymous):

perfect!

OpenStudy (freckles):

\[p=150,210,510,570,870,930 \\ \text{ remember } p=3 \theta \\ 3 \theta=150,210,510,570,870,930\] last step is divide both sides by 3

OpenStudy (freckles):

by the way this is called the lazy way of writing 6 different equations :p

OpenStudy (freckles):

\[\theta=\frac{150}{3}, \frac{210}{3}, \frac{510}{3}, \frac{570}{3}, \frac{870}{3}, \frac{930}{3}\]

OpenStudy (freckles):

and you do the divisions and you are done

OpenStudy (freckles):

but anyways let's go back to the equation you actually had

OpenStudy (anonymous):

awesome! wow.. wouldn't have thought to do that

OpenStudy (freckles):

\[\cos(3 \theta)=\frac{\sqrt{3}}{2}\] \[\text{ where} 0 < \theta<360\]

OpenStudy (freckles):

same thing I would replace 3theta with some other variable like p and then write the solution interval in terms of p as well so I know which p to find

OpenStudy (anonymous):

okay, so 30, and... 330?

OpenStudy (freckles):

right

OpenStudy (freckles):

we will have 2 more rotations of solutions to look at

OpenStudy (freckles):

30 330 30+360 330+360 30+2(360) 330+2(360)

OpenStudy (freckles):

now these will be the solutions for p

OpenStudy (freckles):

and you have p=3 theta so just divide your solutions for p by 3 and you are done

OpenStudy (anonymous):

Wow, thank you so much. I've never been shown that way before

OpenStudy (freckles):

well I hope it is easier for you than the way you have been shown curious what way you have been shown though

OpenStudy (anonymous):

For some, they used factoring. Which confused me more

OpenStudy (freckles):

I don't how you would do factoring for that one... but I do see how one might want to factor: \[\cos^2(x)-2\cos(x)-3=0 \\ \text{ factoring left hand side gives } \\ (\cos(x)-3)(\cos(x)+1)=0\] set both factors equal to 0 you know since a*b=0 then this means a=0 or b=0

OpenStudy (freckles):

\[\cos(x)-3=0 \text{ or } \cos(x)+1=0 \\ \cos(x)=3 \text{ or } \cos(x)=-1 \\ \text{ no solutions for } \cos(x)=3 \\ \text{ but you will have solutions for } \cos(x)=-1 \]

OpenStudy (anonymous):

Yeah it was for this one I had \[2\cos^2 \theta-1=0\]

OpenStudy (freckles):

actually let me try something different for the one you said they factored on... \[\cos(3x)=\frac{\sqrt{3}}{2} \\ \cos(3x)=\cos(2x+x)=\cos(2x)\cos(x)-\sin(2x)\sin(x) \\ \cos(3x)=(\cos^2(x)-\sin^2(x))\cos(x)-(2 \sin(x) \cos(x)) \sin(x) \\ \cos(3x)= \cos^3(x)-\sin^2(x) \cos(x)-2 \sin^2(x) \cos(x) \\ \cos(3x)=\cos^3(x)-(1-\cos^2(x)) \cos(x)-2 (1-\cos^2(x)) \cos(x) \\ \cos(3x)=\cos^3(x)-\cos(x)+\cos^3(x)-2 \cos(x)+2 \cos^3(x) \\ \cos(3x)=4 \cos^3(x)-3 \cos(x) \\ \text{ so you have the equation } \\ 4 \cos^3(x)-3\cos(x)=\frac{\sqrt{3}}{2}\] that is a monster

OpenStudy (freckles):

oh they factored on 2 cos^2(x)-1=0 you can factor on that one or just solve for cos^2(x) first then take square root of both sides either way

OpenStudy (freckles):

if the factoring bit confused with the trig function replace the trig function with some variable as mentioned before

OpenStudy (anonymous):

Holy moly, yeah just looks like a lot more work. Oh true, thats a good idea

OpenStudy (freckles):

\[2p^2-1 =(\sqrt{2} p)^2-(1)^2=(\sqrt{2} p-1)(\sqrt{2}p+1)\]

OpenStudy (freckles):

\[2 \cos^2(x)-1=(\sqrt{2} \cos(x)-1)(\sqrt{2} \cos(x)+1)\]

OpenStudy (freckles):

anyways I have to leave for now

OpenStudy (freckles):

make a new post if you need anymore help

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (freckles):

np :)